OpenStudy (anonymous):

find the values of x and y that maximize or minimize the objective function for each of these graphs.

OpenStudy (anonymous):

heres the link that has the graphs

OpenStudy (anonymous):

its the second to last question. the one with the graphs

OpenStudy (anonymous):

@jordanloveangel

OpenStudy (jordanloveangel):

x + y < 8 2x + y < 10 x > 0, y > 10 (0, 5); maximum value is 100 (1, 7); maximum value is 220 (2, 6); maximum value is 280 (5, 0); maximum value is 400

OpenStudy (jordanloveangel):

You did it. x=5 y=0 How you do it is graph the lines on a piece of graph paper. The lines going less then shade the lines below the line. The lines going greater than shade above the area. Now you will see a shape that is formed that inside it every shading is in it. Next what you do is find the points that make the shape. Then plug in the x's and y's for each point into the original equation in which you did not write down. The highest number you get is maximizing. The lowest is minimizing.

OpenStudy (anonymous):

but what about the second one?

OpenStudy (anonymous):

honestly, my friend helped me with the first one but I dont know how she did it

OpenStudy (jordanloveangel):

You have all the answers at your disposal. At (0,5), it is minimum at 100 and at (5,0), it is maximum at 400. this is for the first one and one sec m i try 2 do the 2end one

OpenStudy (anonymous):

okay

OpenStudy (jordanloveangel):

idk the 2end one

OpenStudy (anonymous):

awh okay. do you know anyone else that might be able to help?

OpenStudy (anonymous):

@jordanloveangel ?

OpenStudy (jordanloveangel):

1 sec

OpenStudy (jordanloveangel):

@magan @MarziaJackson$ @BunnyBree @Tazmaniadevil @taylor12344 @Walkerman15 @waterineyes @Whitemonsterbunny17 @eliciagarza605273 @emo__gamer101 @redeemed1312 @Rockiee @THIS,KILLCHIC @Tina_Russo @young.money @itsbribro @iRespect @priecetheveil @alesia/bailey @Angelina_Vella @AJLee5986 @secret66 @sailawaywithme @sandrayanez @Savane0621 @savannah152426 @stellawalkerson @SlowerThePriest @dan815 @danish071996 @daylon @DullJackel09 @Gabylovesyou @Here_to_Help15 @Jamielynne @JaScOk @jordanloveangel @Jhannybean @KaMaddie @KyanTheDoodle @KailaMorgan7777 @lakeykey @love_to_love_you @limelife @ZoMbAsH @Countrygirl23 @Clueless156 @Crissy15 @cheerleading4life<3 @confusegirllll @CoolKatt16 @chrissy_xoxo @Clalgee @ValarieAnne @BabyGirlAj @baking4lyf @Broken_heartedgirl123 @BunnyBree @Natalia.lovs.you @morganchavez @marylou004 @Mesopretty @myininaya @Mimi_x3 @mkerr017 @mcnabkt @...Dimps...

OpenStudy (alesia/bailey):

2x+y=0 x-3y=-7 so you choose one of the problems, ill pick 2x+y=0 your first goal is to isolate one of the variables, x or y ill do y 2x+y=0 -2x -2x y=0-2x so you plug that into your other problem x-3y=-7 x-3(0-2x)=-7 you plugged in your y so now your trying to figure out x x-0-6x=-7 combine the x's 0-7x=-7 -0 -0 -7x=-7 /-7 /-7 x=1 so now you know what x is now you plug it back in with your first problem that you isolated y=0-2x you fill in x with 1, because x is 1 y=0-2(1) y=0-2 y=-2 y=-2, x=1 from there you can plug it in with your graph. (I might've messed up with through the problem getting the wrong answer, but that how you solve.